Car A and car B both travel 560 miles. Car A drove at an average speed of 10 miles per hour faster than car B and made the trip in one hour less time. Find the average speeds of both cars.
Answer by palanisamy(496)

Let the speed of car A be x.
Then,the speed of car B is x-10.
Distance travelled = 560 mile.
Time taken by car A = 560/x
Time taken by car B = 560/(x-10)
Given
560/x -10 = 560/(x-10)
(5600-10x)/10x = 560/(x-10)
(5600-10x)(x-10) = 5600x
5600x-56000-10x^2+100x = 5600x
-10x^2+100x-56000 = 0
x^2-10x+5600 = 0
(x-80)(x+70) = 0
x = 80 or -70
x cannot be negative.
So, x = 80
Therefore,the speed of car A is 80 miles/h
and,the speed of car B is x-10 = 80-10 = 70 miles/h